40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 280     Entry time: Fri Jan 27 14:56:55 2017
Author: Gabriele 
Type: Optics 
Category: Daily Progress 
Subject: Optical lever setup completed 

The four optical levers are completely installed and aligned to a horizontal reference.

ELOG V3.1.3-